a fenced yard with a lenght of 110 feet and width of 65 feet.
Thanks Tutor.

Answer by tutorcecilia(2152)

Are you asking for the perimeter or the area?
Area = (Length)(Width)
A=(110)(65)=7510 sq. feet
.
Perimeter=2(Length + width)
P=2(110+65)
P=2(175)
P=350 feet